Calculate the acceleration and the distance covered when train initially travelling at a speed of 10m/s, increases its speed to 25m/s in 20s
1
Expert's answer
2020-01-06T10:17:27-0500

The acceleration of the train


"a=\frac{v_f-v_i}{t}=\frac{25-10}{20}=0.75\:\rm m/s^2"

The distance traveled by train


"d=\frac{v_i+v_f}{2}t=\frac{10+25}{2}\times 20=350\:\rm m"